# Self-Routing RAG: Binding Selective Retrieval with Knowledge Verbalization

### v2

@@ C.2 Discussion: Further Applications
Querying different knowledge sources brings distinct latency cost. SR-RAG’s preference label can be constructed such that the cheapest useful knowledge source is preferred, improving the latency while preventing the performance degradation. In addition, varied levels of compute are required to verbalize the required knowledge for different types of queries. SR-RAG can be used to model this pattern by formulating multiple internal knowledge sources associated with different levels of compute. This philosophy connects with the literature on adaptive reasoning budget (Lou et al., 2025).
@@
Appendix D Further Analyses
D.1 Self-Routing RAG: Hyperparameters
@@
In SR-RAG, the default design for source preference labeling is collecting the knowledge from each source and selecting the top-ranked source in top 50% knowledge contexts in terms of contribution to the likelihood of the answer. In this section, we compare this strategy to a number of alternative heuristics:
@@
In Table˜7, we present the F1 score and AUROC using Llama-2-7B-Chat as the LLM. Correctness is defined as preferring self-verbalized knowledge when it is better than or equal to performing retrieval in terms of the downstream question answering performance. Compared to three baseline methods, the heuristics SR-RAG uses achieves the best F1 and AUROC over the short-form and closed-set subset of the training set. Building upon this paper’s results, future work can further study leveraging more advanced uncertainty quantification methods (Wu et al., 2024b; Moskvoretskii et al., 2025b) to build more accurate source selection labels.
@@
Table 7: F1 and AUROC scores for different source labeling methods across datasets. Llama-2-7B-Chat is used as the LLM. The best score per column is boldfaced.
@@
We further study the impact of layer selection on SR-RAG’s performance in Figure˜6 part 3 (green). Overall, we observe that using a middle layer’s hidden states has more superior performance over using the last layer’s. In the next section, we visualize the hidden state space across LLMs and show that middle layers tend to learn better representations for the model uncertainty as well as task types.
D.2 Hidden State Space of the Source Reflection Token
Does SR-RAG training allow the model to condense their knowledge of task, model uncertainty, and source characteristics into the representation of <EOQ>?
@@ D.2 Hidden State Space of the Source Reflection Token
Table 8: Comparison between training with self-generated knowledge versus knowledge distillation for the verbalization branch in SR-RAG. Llama-2-7B-Chat is fine-tuned either with self-verbalized knowledge or the knowledge verbalized from Llama-3.3-70B-Instruct.
D.3 Can SR-RAG learn more from self-generated knowledge?
@@
D.4 Qualitative Study
In Figure˜7 and Figure˜8, we show two qualitative examples from TriviaQA with the fine-tuned Llama-2-7B-Chat as the backbone LLM. For both examples, we show the LLM’s prediction conditioned on both retrieved knowledge and the verbalized knowledge. While the verbalized knowledge is more concise than retrieved paragraphs, it better identifies the correct answers for both questions. This is because the core information (i.e., "first name", "island") only takes up a small portion of the query. As a result, dense retrieval is noise-prone, bringing in only distantly relevant information. By contrast, the LLM as a knowledge generator dynamically adjusts its focus, guiding the knowledge generation to include the core information. As a result, the LLM can correctly answer both questions with self-verbalized knowledge but provides incorrect answers with the low-quality retrieved knowledge. In the source selection process, the LLM directly predicts the source preference $p_{M}(\text{{\textless s\textgreater}}|q)$. In addition, questions from the train set sharing the same topic and difficulty level are retrieved to form the source preference distribution $p_{D}(\text{{\textless s\textgreater}}|q)$. In the first example (Figure˜7), both distributions prefer the LLM itself as the knowledge source. By contrast, in the second example (Figure˜8), $p_{M}(\text{{\textless s\textgreater}}|q)$ has a high uncertainty and $p_{D}(\text{{\textless s\textgreater}}|q)$ corrects the source distribution to select the LLM itself as the knowledge source.
@@ D.4 Qualitative Study
Figure 7: Example of SR-RAG inference where $p_{D}(\text{{\textless s\textgreater}}|q)$ agrees with $p_{M}(\text{{\textless s\textgreater}}|q)$. As the model itself is selected as the knowledge source, the retrieval will be abstained. The retrieved knowledge and the corresponding prediction are shown only for illustration.
@@
Figure 8: Example of SR-RAG inference where $p_{D}(\text{{\textless s\textgreater}}|q)$ corrects the source selection from $p_{M}(\text{{\textless s\textgreater}}|q)$. The model itself is selected as the knowledge source.
Figure 9: Hidden states of <EOQ> across three LLMs on PopQA, visualized using t-SNE. Llama = Llama-2-7B-Chat, Phi = Phi-3.5-Mini-Instruct, and Qwen = Qwen2.5-7B-Instruct.
Figure 10: Hidden states of <EOQ> across three LLMs on a random sample of 500 data points in the SR-RAG training set, visualized using t-SNE. Llama = Llama-2-7B-Chat, Phi = Phi-3.5-Mini-Instruct, and Qwen = Qwen2.5-7B-Instruct.

### v3

@@ B.2 Discussion: Further Applications
Querying different knowledge sources brings distinct latency costs. SR-RAG’s preference label can be constructed such that the cheapest useful knowledge source is preferred, improving the latency while preventing performance degradation. In addition, varied levels of compute are required to verbalize the required knowledge for different types of queries. SR-RAG can be used to model this pattern by formulating multiple internal knowledge sources associated with different levels of compute. This philosophy connects with the literature on adaptive reasoning budget (lou2025adacot).
@@
Appendix C Further Analyses
C.1 Self-Routing RAG: Hyperparameters
@@
In SR-RAG, the default design for source preference labeling is to collect the knowledge from each source and to select the top-ranked source among the top 50% of knowledge contexts in terms of contribution to the likelihood of the answer. In this section, we compare this strategy to a number of alternative heuristics:
@@
In Table˜7, we present the F1 score and AUROC using Llama-2-7B-Chat as the LLM. Correctness is defined as preferring self-verbalized knowledge when it is better than or equal to performing retrieval in terms of the downstream question answering performance. Compared to three baseline methods, the heuristics SR-RAG uses achieves the best F1 and AUROC over the short-form and closed-set subset of the training set. Building upon this paper’s results, future work can further study leveraging more advanced uncertainty quantification methods (wu-etal-2024-synchronous; DBLP:journals/corr/abs-2501-12835) to build more accurate source selection labels.
@@
Table 7: Source labeling accuracy of different heuristics. F1 and AUROC scores for different methods across datasets. Llama-2-7B-Chat is used as the LLM. The best score per column is boldfaced.
@@
We further study the impact of layer selection on SR-RAG’s performance in Figure˜6 part 3 (green). Overall, we observe that using a middle layer’s hidden states has better performance over using the last layer’s. In the next section, we visualize the hidden state space across LLMs and show that middle layers tend to learn better representations for the model uncertainty as well as task types.
C.2 Hidden State Space of the Source Reflection Token
Does SR-RAG training allow the model to condense its knowledge of tasks, model uncertainty, and source characteristics into the representation of <EOQ>?
@@ C.2 Hidden State Space of the Source Reflection Token
Table 8: Effect of knowledge distillation. Comparison between training with self-generated knowledge versus knowledge distillation for the verbalization branch in SR-RAG. Llama-2-7B-Chat is fine-tuned either with self-verbalized knowledge or the knowledge verbalized from Llama-3.3-70B-Instruct.
C.3 Can SR-RAG learn more from self-generated knowledge?
@@
C.4 Discussion: Efficiency and Applicability
In this section, we briefly discuss the efficiency and applicability of SR-RAG.
Training Efficiency
SR-RAG introduces additional offline computation because source labels are constructed by collecting multiple candidate contexts per source (retrieved passages and self-verbalizations) and scoring them with answer likelihood (Section˜A.2). We view this cost as a one-time post-training expense that can be amortized across deployments, and it requires no extra human annotation or stronger teacher models. In addition, our two-stage recipe mirrors common modern post-training pipelines that combine supervised fine-tuning with preference alignment, and it can be incorporated without substantial changes to standard training infrastructure.
Inference Efficiency
SR-RAG preserves efficient inference by executing routing, knowledge collection, and answer generation in a single left-to-right pass (Figure˜1). When the internal source is selected, the model produces one compact knowledge verbalization instead of running multi-sample elicitation; when an external source is selected, decoding pauses only to issue retrieval, and overall latency is largely governed by retrieval frequency (Figures˜4 and A.3). The kNN policy adds a lightweight computation when the datastore is cached on GPU, and our hyperparameter analyses show that routing quality is not highly sensitive to datastore size, suggesting that compact or compressed datastores can be effective (Figure˜6).
Applicability
SR-RAG is designed for practical routing settings where systems must choose among heterogeneous sources. First, SR-RAG achieves favorable accuracy-efficiency trade-offs without dataset-specific threshold tuning, which reduces the manual effort required to deploy selective retrieval across diverse workloads (Figure˜4). Second, the policy datastore improves interpretability because it stores explicit source assignments that can be inspected and debugged, rather than relying solely on opaque likelihood thresholds. Third, the same datastore also enables controllability: practitioners can audit and modify routing behavior by editing, filtering, or augmenting stored examples (for instance, enforcing least-privilege routing or overriding problematic regions of the decision boundary). Finally, SR-RAG naturally extends beyond the two-source setup, and our three-source experiments illustrate that the framework can route to specialized corpora (e.g., PubMed) when needed while retaining strong performance on general workloads (Appendix˜B).
C.5 Qualitative Study
In Figure˜7 and Figure˜8, we show two qualitative examples from TriviaQA with the fine-tuned Llama-2-7B-Chat as the backbone LLM. For both examples, we show the LLM’s prediction conditioned on both retrieved knowledge and the verbalized knowledge. While the verbalized knowledge is more concise than retrieved paragraphs, it better identifies the correct answers for both questions. This is because the core information (i.e., "first name", "island") only takes up a small portion of the query. As a result, dense retrieval is noise-prone and brings in only distantly relevant information. By contrast, the LLM as a knowledge generator dynamically adjusts its focus, guiding the knowledge generation to include the core information. As a result, the LLM can correctly answer both questions with self-verbalized knowledge but provides incorrect answers with the low-quality retrieved knowledge. In the source selection process, the LLM directly predicts the source preference $p_{M}(\text{{\textless s\textgreater}}|q)$. In addition, questions from the train set sharing the same topic and difficulty level are retrieved to form the source preference distribution $p_{D}(\text{{\textless s\textgreater}}|q)$. In the first example (Figure˜7), both distributions prefer the LLM itself as the knowledge source. By contrast, in the second example (Figure˜8), $p_{M}(\text{{\textless s\textgreater}}|q)$ has a high uncertainty and $p_{D}(\text{{\textless s\textgreater}}|q)$ corrects the source distribution to select the LLM itself as the knowledge source.
@@ C.5 Qualitative Study
Figure 7: Qualitative Example 1. Example of SR-RAG inference where $p_{D}(\text{{\textless s\textgreater}}|q)$ agrees with $p_{M}(\text{{\textless s\textgreater}}|q)$. As the model itself is selected as the knowledge source, the retrieval will be abstained. The retrieved knowledge and the corresponding prediction are shown only for illustration.
@@
Figure 8: Qualitative Example 2. Example of SR-RAG inference where $p_{D}(\text{{\textless s\textgreater}}|q)$ corrects the source selection from $p_{M}(\text{{\textless s\textgreater}}|q)$. The model itself is selected as the knowledge source.
Figure 9: Hidden State Visualization 1. Hidden states of <EOQ> across three LLMs on PopQA, visualized using t-SNE. Llama = Llama-2-7B-Chat, Phi = Phi-3.5-Mini-Instruct, and Qwen = Qwen2.5-7B-Instruct.
Figure 10: Hidden State Visualization 2. Hidden states of <EOQ> across three LLMs on a random sample of 500 data points in the SR-RAG training set, visualized using t-SNE. Llama = Llama-2-7B-Chat, Phi = Phi-3.5-Mini-Instruct, and Qwen = Qwen2.5-7B-Instruct.